OpenStudy (anonymous):

+ & - Radical Expressions help please!!! (give me a moment to insert the equation)

OpenStudy (anonymous):

\[\LARGE 4\sqrt{2} + 7\sqrt{2} - 3\sqrt{2}\] ANSWER CHOICES: A) \[\Large 2\sqrt{8}\] B) \[\Large 8\sqrt{2}\] C) \[\Large 8\sqrt{6}\] D) \[\Large 6\sqrt{8}\]

OpenStudy (anonymous):

Can somebody please tell me how to do this?

OpenStudy (anonymous):

\[4x+7x-3x=?\]

OpenStudy (anonymous):

just the same as the algebra question "combine like terms"

OpenStudy (anonymous):

I still don't understand...

OpenStudy (anonymous):

@jim_thompson5910 Can you help me?? :)

jimthompson5910 (jim_thompson5910):

What do you get when you add 2x to 3x (ie how how do you combine these like terms)?

OpenStudy (anonymous):

5x?

jimthompson5910 (jim_thompson5910):

how did you do that

OpenStudy (anonymous):

I added the 2 & 3 & then put the x back on.

jimthompson5910 (jim_thompson5910):

good, that's exactly what satellite73 is saying and exactly what you do so for instance, if you wanted to add \[\Large 5\sqrt{3}+7\sqrt{3}\], you just add the numbers in front of the square root of 3 (ignore the square root of 3) to get 5+7 = 12 Then you tack the square root of 3 back on the result you just got to say \[\Large 5\sqrt{3}+7\sqrt{3}=12\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

So what do you get when you simplify \[\Large 4\sqrt{2} + 7\sqrt{2} - 3\sqrt{2}\]

OpenStudy (anonymous):

\[\LARGE 8\sqrt{2}\] ?? :)

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

that's all there is to it

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 \[\Large (12\sqrt{5} - 7\sqrt{7}) + (\sqrt{5} - 7\sqrt{7})\] I tried to do the same thing with this one but I got confused... Help, please?

jimthompson5910 (jim_thompson5910):

well this time, you can only do it for the like terms

OpenStudy (anonymous):

\[\Large (12\sqrt{5} - 7\sqrt{7}) + (6\sqrt{5} - 7\sqrt{7})\]

OpenStudy (anonymous):

Corrected^

jimthompson5910 (jim_thompson5910):

for instance, you cannot add the numbers in front of the sqrt(5) and sqrt(7)

jimthompson5910 (jim_thompson5910):

the numbers in front of the sqrt(5) are...???

OpenStudy (anonymous):

12 & 6

OpenStudy (anonymous):

Would I add them?

jimthompson5910 (jim_thompson5910):

so they add to 18, which means that the final result for the sqrt(5) term is \[\Large 12\sqrt{5}+6\sqrt{5} = 18\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

yes you would

OpenStudy (anonymous):

And so it would be 18sqrrt5 & that's it b/c the 7's cancel each other out?

OpenStudy (anonymous):

Cos they equal zero?

jimthompson5910 (jim_thompson5910):

no, not quite

jimthompson5910 (jim_thompson5910):

-7 + -7 isn't 0

OpenStudy (anonymous):

oh... it's -14

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so \[\LARGE 18\sqrt{5} and -14\sqrt{7}\] ??

jimthompson5910 (jim_thompson5910):

then you just write them as one expression, so the answer would be \[\Large 18\sqrt{5} -14\sqrt{7}\] you can't combine these because they are not like terms.

OpenStudy (anonymous):

oh, okay! Thank you :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Would \[\Large \sqrt{10} \times \sqrt{8}\] be \[\Large \sqrt{80}\] ??

jimthompson5910 (jim_thompson5910):

yes and you can simplify that

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Wait... to \[\Large 4\sqrt{5} \] ?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

do you know how to get that?

OpenStudy (anonymous):

The same way you taught me to simplify the other day :)

jimthompson5910 (jim_thompson5910):

ok great, just checking

OpenStudy (anonymous):

Whoah! How do I do this one? \[\LARGE \sqrt{5}(8 + 3\sqrt{6})\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

treat sqrt(5) just like any number and distribute it through

OpenStudy (anonymous):

so 8*5 is 40 5*3=15 5*6 = 30

jimthompson5910 (jim_thompson5910):

kinda, but you multiply sqrt(5) by 8 and sqrt(5) by 3*sqrt(6)

OpenStudy (anonymous):

I'm confused..

jimthompson5910 (jim_thompson5910):

sqrt(5) times 8 = ??

OpenStudy (anonymous):

... 17-ish?? Idk.

jimthompson5910 (jim_thompson5910):

no, it's just 8*sqrt(5)

OpenStudy (anonymous):

oh.... okay ha

jimthompson5910 (jim_thompson5910):

sqrt(5) times 3*sqrt(6) = ???

OpenStudy (anonymous):

Idk...